# Diederich–Fornæss index and global regularity of the complex Green operator: domains with comparable Levi eigenvalues

### v1

@@ 1. Introduction
$DF(\Omega)$ refers to the Diederich–Fornæss index of $\Omega$, and $D(q_{0})$ is a comparability condition bearing on sums of $q_{0}$ eigenvalues of the Levi form: precise definitions for both, as well as for the related condition $\widetilde{D}(q)$, are in Section 2. Observe that $\min\{q_{0},n-1-q_{0}\}$ or $\max\{q_{0},n-1-q_{0}\}$ equals $q_{0}$, depending on whether $q_{0}\leq(n-1)/2$ or $q_{0}\geq(n-1)/2$, respectively. In particular, the range of allowable $q$’s is symmetric about (n1)/2(n-1)/2, in line with the expectation that estimates on the boundary should hold at symmetric form levels simultaneously (see the discussion around (5.40)). The estimates in Theorem 1 are not independent: (3) holds for (0,q)(0,q)–forms if and only if (2) holds at levels (q1)(q-1), $q$, and $(q+1)$ ([18]), and (1) holds if and only if (2) holds (appropriately formulated when $q=0$ or q=(n1)q=(n-1), [35]).
@@ 3. DF–index, commutators, and D’Angelo forms
Moreover, there is $\delta_{\eta}>0$ such that this estimate holds uniformly on the level sets $\{\rho_{\eta}=-\delta\}$ for $\delta<\delta_{\eta}$. Roughly speaking, the Hessian of $h_{\eta}$ appears because $\overline{\partial}\alpha_{\rho_{\eta}}$ in (3.5) equals $\overline{\partial}(dh_{\eta}+\alpha_{\rho})$ (some care is needed though; see [10], Section 4.3, especially Lemmas 4.6 and 4.7, for details), and the $M_{\eta}$–term is needed for when $L(p)\notin\mathcal{N}_{p}$ (see [28], section 3). In order to obtain an estimate for forms from (3.7), we define auxiliary vector fields associated to forms as follows. For a (0,q)(0,q)–form $u=\sideset{}{{}^{\prime}}{\sum}_{|K|=q}u_{K}d\overline{z_{K}}$ and an increasing (q1)(q-1)–tuple $J$, set $L^{J}_{u}=\sum_{j}u_{jJ}\partial/\partial z_{j}$ (when $q=1$, we understand $J$ to be the empty tuple, and $L_{u}=\sum_{j}u_{j}\partial/\partial z_{j}$). When $u$ is a form on the boundary, then for all $J$, $L^{J}_{u}\rho=\sum_{j}u_{jJ}(\partial\rho/\partial z_{j})=0$, so that (3.7) does apply to $L^{J}_{u}$. Note that $L_{u}^{J}$ is not invariant under a change of frame from Euclidean to a special boundary frame (unless $q=1$). However, the quantity that will matter, namely $\sum^{{}^{\prime}}_{J}\int_{b\Omega}|\alpha_{\rho_{\eta}}(L_{u}^{J})|^{2}$ (the quantity on the left hand side of (3.8) below) is, up to a constant factor. The reason is that each $L_{u}^{J}$ is a linear combination, with smooth coefficients independent of $u$, of the corresponding vectors in the other frame (this observation from [28], page 5417, is based on the transformation formulas between the two orthonormal frames). Integrating (3.7) for $L_{u}^{J}$ over the boundary and summing over $J$ gives
(3.8) $\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}|\alpha_{\rho_{\eta}}(L^{J}_{u})|^{2}\leq C(1-\eta)\left(\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}\sum_{j,k}\frac{\partial^{2}(-h_{\eta})}{\partial z_{j}\partial\overline{z_{k}}}u_{jJ}\overline{u_{kJ}}+2\|u\|^{2}\right)\\ +M_{\eta}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}\sum_{j,k}\frac{\partial^{2}\rho}{\partial z_{j}\partial\overline{z_{k}}}u_{jJ}\overline{u_{kJ}}\;\;.$
Moreover, there is $\delta_{\eta}$ so that this estimate holds uniformly over the level sets $\rho_{\eta}=-\delta$ for $0\leq\delta<\delta_{\eta}$. This estimate is the starting point, in the next section, for deriving the crucial $L^{2}$–type estimate needed in the proof of Theorem 1.
@@ 4. The main estimate
The estimate in Proposition 1 is, for our purposes, the principal consequence of having DF–index 1.
@@
(4.1) $\sideset{}{{}^{\prime}}{\sum}_{|J|=q-1}\int_{b\Omega}\left|\alpha_{\rho_{\eta}}(L_{u}^{J})\right|^{2}\leq C(1-\eta)\left(\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\right)+C_{\eta}\|u\|_{-1}^{2}\,$,
@@
It suffices to prove the proposition for forms supported in suitable small patches of the boundary (independent of $\eta$), the general case then follows via a partition of unity. For $u$ supported in a special boundary chart as in Proposition 2 in section 6, we look at the term involving the Hessian of $h_{\eta}$ on the right hand side of (3.8). Proposition 2, with $\varphi=-h_{\eta}$, and $(e^{-h_{\eta}}/2)u$ in place of $u$, gives
[…]
here $\overline{L}(\cdot)$ denotes the vector of all $\overline{L_{j}}$–derivatives, $u_{j}$ are the components of $u$ in the special boundary chart, $\widetilde{u_{J}}$ denotes the Euclidean components of $u$, $c_{jk}$ are the coefficients of the Levi form in the basis given by the boundary frame, and the constant in the $\mathcal{O}(\|u\|^{2})$ is independent of $\eta$. Since $T$ is purely imaginary and $h_{\eta}$ is real, we get $Th_{\eta}$ is purely imaginary and
(4.3) $\text{Re}\sum_{j,k}c_{jk}(Th_{\eta})u_{jK}\overline{u_{kK}}=\text{Re}\left((Th_{\eta})\sum_{j,k}c_{jk}u_{jK}\overline{u_{kK}}\right)=0\;,$
since pseudoconvexity implies that the summation above results in a (non-negative) real number. We rewrite (4.2) as
(4.4) $\begin{split}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}(-h_{\eta})}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}&\lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}\\ &\quad-2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\big((T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})u_{jK}\big)\overline{u_{kK}}+\mathcal{O}(\left\|u\right\|^{2}).\end{split}$
@@
(4.5) $\begin{split}\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(L_{u}^{K})\right|^{2}&\lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}\\ &\qquad-2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\big((T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})u_{jK}\big)\overline{u_{kK}}+\mathcal{O}(\left\|u\right\|^{2})\Big)\\ &\qquad\qquad+M_{\eta}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}u_{jK}\overline{u_{kK}}.\end{split}$
@@
(4.6) $\begin{split}\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(L_{u}^{K})\right|^{2}&\lesssim(1-\eta)\bigg(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}\\ &\qquad-2\text{Re}\big(\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{jk}\int_{b\Omega}c_{jk}\big((T-\tilde{M}_{\eta})u_{jK}\big)\overline{u_{kK}}\,\big)+\mathcal{O}(\left\|u\right\|^{2})\bigg)\end{split}$
@@
It now suffices to prove that $\sum^{{}^{\prime}}_{|J|=q-1}\int_{b\Omega}|\alpha_{\rho_{\eta}}(L_{\mathcal{P}^{k,\eta}u}^{J})|^{2}$ is dominated by the right hand side of (4.1), for $k=+,0,$-.
@@
(4.22) $\sideset{}{{}^{\prime}}{\sum}_{J}|\alpha_{\rho_{\eta}}(L^{J}_{\mathcal{P}^{+,\eta}u})|^{2}\\ \lesssim(1-\eta)\left(\|\overline{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b,-h_{\eta}}^{*}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\right)\\ +C_{\eta}\|u\|_{-1}^{2}\;.\;\;\;\;\;\;\;\;$
@@
(4.23) $2\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}(-h_{\eta})}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}\\ +2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}(T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu}-\frac{1}{2}Th_{\eta})u_{jK}\overline{u_{kK}}+3q\left\|\overline{L}u\right\|^{2}\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\mathcal{O}(\left\|u\right\|^{2})+3q\left\|\overline{L}u\right\|^{2}.$
@@
(4.25) $2\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}(-h_{\eta})}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}+2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}(T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})u_{jK}\overline{u_{kK}}\\ -3q\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}(c_{jj}Tu_{J})\overline{u_{J}}+3q\left\|Lu\right\|^{2}+\mathcal{O}(\left\|{L}u\right\|\left\|u\right\|)+\mathcal{O}(\left\|\overline{L}u\right\|\left\|u\right\|)\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\mathcal{O}(\left\|u\right\|^{2})+3\left\|\overline{L}u\right\|^{2}.$
We have also used (4.3). Using $\|\overline{L}u\|^{2}\geq 0$, the Cauchy-Schwarz inequality, maximal estimates, and (2.2) gives
(4.26) $2\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}(-h_{\eta})}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}\\ +2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}(T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})u_{jK}\overline{u_{kK}}-3q\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}(c_{jj}Tu_{J})\overline{u_{J}}\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}.$
@@
(4.27) $2\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\dfrac{\partial^{2}(-h_{\eta})}{\partial z_{j}\partial\overline{z}_{k}}\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}\\ +\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\left(\left(2c_{jk}-3\delta_{jk}\text{Tr}(c)\right)T+\frac{\partial h_{\eta}}{\partial\nu}c_{jk}\right)u_{jK}\overline{u_{kK}}\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}\;,$
where $Tr(c)$ denotes the trace of the Levi form. Multiplying both sides of (4.27) by $(1-\eta)$, then adding $M_{\eta}\sum_{K}^{\prime}\sum_{j,k}\int_{b\Omega}({\partial^{2}\rho}/{\partial z_{j}\partial\overline{z}_{k}})\widetilde{u_{jK}}\overline{\widetilde{u_{kK}}}=M_{\eta}\sum^{{}^{\prime}}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}u_{jK}\overline{u_{kK}}$, and invoking (3.8) and (2.2) leads to
(4.28) $\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}|\alpha_{\rho_{\eta}}(L_{u}^{J})|^{2}\\ \lesssim(1-\eta)\bigg(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}\\ -\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\left(2c_{jk}-3\delta_{jk}\text{Tr}(c)\right)Tu_{jK}\overline{u_{kK}}+\tilde{M}_{\eta}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}u_{jK}\overline{u_{kK}}\bigg),$
where $\tilde{M}_{\eta}=\frac{M_{\eta}}{(1-\eta)}$.
We now need an upper bound on the last two terms on the right hand side of (4.28), with $u$ replaced by $\mathcal{P}^{-,\eta}u$, equivalently, a lower bound on

### v2

@@ 1. Introduction
$DF(\Omega)$ refers to the Diederich–Fornæss index of $\Omega$, and $D(q_{0})$ is a comparability condition bearing on sums of $q_{0}$ eigenvalues of the Levi form introduced by Derridj ([12]) (see also Koenig ([20]): precise definitions for both, as well as for the related condition $\widetilde{D}(q)$, are in Section 2. Observe that $\min\{q_{0},n-1-q_{0}\}$ or $\max\{q_{0},n-1-q_{0}\}$ equals $q_{0}$, depending on whether $q_{0}\leq(n-1)/2$ or $q_{0}\geq(n-1)/2$, respectively. In particular, the range of allowable $q$’s is symmetric about (n1)/2(n-1)/2, in line with the expectation that estimates on the boundary should hold at symmetric form levels simultaneously (see the discussion around (5.40)). The estimates in Theorem 1 are not independent: (3) holds for (0,q)(0,q)–forms if and only if (2) holds at levels (q1)(q-1), $q$, and $(q+1)$ ([18]), and (1) holds if and only if (2) holds (appropriately formulated when $q=0$ or q=(n1)q=(n-1), [35]).
@@ 3. DF–index, commutators, and D’Angelo forms
Moreover, there is $\delta_{\eta}>0$ such that this estimate holds uniformly on the level sets $\{\rho_{\eta}=-\delta\}$ for $\delta<\delta_{\eta}$. Roughly speaking, the Hessian of $h_{\eta}$ appears because $\overline{\partial}\alpha_{\rho_{\eta}}$ in (3.5) equals $\overline{\partial}(dh_{\eta}+\alpha_{\rho})$ (some care is needed though; see [10], Section 4.3, especially Lemmas 4.6 and 4.7, for details), and the $M_{\eta}$–term is needed for when $L(p)\notin\mathcal{N}_{p}$ (see [28], section 3). In order to obtain an estimate for forms from (3.7), we define auxiliary vector fields associated to forms as follows. For a (0,q)(0,q)–form $u=\sideset{}{{}^{\prime}}{\sum}_{|K|=q}u_{K}d\overline{z_{K}}$ and an increasing (q1)(q-1)–tuple $J$, set $L^{J}_{u}=\sum_{j}u_{jJ}\partial/\partial z_{j}$ (when $q=1$, we understand $J$ to be the empty tuple, and $L_{u}=\sum_{j}u_{j}\partial/\partial z_{j}$). When $u$ is a form on the boundary, then for all $J$, $L^{J}_{u}\rho=\sum_{j}u_{jJ}(\partial\rho/\partial z_{j})=0$, so that (3.7) does apply to $L^{J}_{u}$. In sections 4 and 5, we will work in special boundary charts, with the usual (local) orthonormal basis $L_{1},L_{2},\cdots,L_{n}$ for vetcor fields of type $(1,0)$, and dual basis $\omega_{1},\cdots,\omega_{n}$. If $u=\sum^{{}^{\prime}}u_{K}\overline{\omega_{K}}$, the analogous construction in the boundary chart is $\tilde{L}_{u}^{J}=\sum_{j}u_{jJ}L_{j}$. Applying (3.7) to $\tilde{L}_{u}^{J}$, and summing over $J$, gives
(3.8) $\sideset{}{{}^{\prime}}{\sum}_{J}|\alpha_{\rho_{\eta}}(\tilde{L}^{J}_{u})|^{2}\leq C(1-\eta)\left(\sideset{}{{}^{\prime}}{\sum}_{J}\partial\overline{\partial}(-h_{\eta})\big(\tilde{L}_{u}^{J}\wedge\overline{\tilde{L}_{u}^{J}}\big)+2\|u\|^{2}\right)\\ +M_{\eta}\sideset{}{{}^{\prime}}{\sum}_{J}\sum_{j,k}c_{jk}u_{jJ}\overline{u_{kJ}}\;\;,$
where $(c_{jk})$ are the coefficients of the Levi form in the boundary chart. There is $\delta_{\eta}$ so that this estimate holds uniformly over the level sets $\rho_{\eta}=-\delta$ for $0\leq\delta<\delta_{\eta}$ (inherited from (3.7)). The above construction is not invariant (unless $q=1$). However, the quantity that will matter, namely $\sum^{{}^{\prime}}_{J}|\alpha_{\rho_{\eta}}(L_{u}^{J})|^{2}$ (respectively $\sum^{{}^{\prime}}_{J}|\alpha_{\rho_{\eta}}(\tilde{L}_{u}^{J})|^{2}$) is, up to a constant factor. The reason is that each $L_{u}^{J}$ is a linear combination, with smooth coefficients independent of $u$, of the $\tilde{L}_{u}^{J}$, and vice versa (this observation from [28], page 5417, is based on the transformation formulas between the two orthonormal frames). Estimate (3.8) is the starting point, in the next section, for deriving the crucial $L^{2}$–type estimate needed in the proof of Theorem 1.
@@ 4. The main estimate
To simplify notation from section 3, now that there is an additional subscript $\eta$, we replace the double subscripts $\rho_{\eta}$ by the simple subscript $\eta$. That is $T_{\rho_{\eta}}$ becomes $T_{\eta}$, $\alpha_{\rho_{\eta}}$ becomes $\alpha_{\eta}$, and so on. The estimate in Proposition 1 is, for our purposes, the principal consequence of having DF–index 1.
@@
(4.1) $\sideset{}{{}^{\prime}}{\sum}_{|J|=q-1}\int_{b\Omega}\left|\alpha_{\eta}(L_{u}^{J})\right|^{2}\leq C(1-\eta)\left(\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\right)+C_{\eta}\|u\|_{-1}^{2}\,$,
@@
It suffices to prove the proposition for forms supported in suitable small patches of the boundary (independent of $\eta$), the general case then follows via a partition of unity. In view of the discussion at the end of the previous section, we may replace the $L_{u}^{J}$ in (4.1) by $\tilde{L}_{u}^{J}$. For $u$ supported in a special boundary chart as in Proposition 2 in section 6, we look at the term involving the Hessian of $h_{\eta}$ on the right hand side of (3.8). Proposition 2, with $\varphi=-h_{\eta}$, and $(e^{-h_{\eta}}/2)u$ in place of $u$, gives
(4.2) $\|\overline{L}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\left(\left(T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu}\right)u_{jK}\right)\overline{u_{kK}}\\ +2\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\partial\overline{\partial}(-h_{\eta})\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)+\mathcal{O}(\left\|u\right\|^{2})\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}\;$;
here $\overline{L}(\cdot)$ denotes the vector of all $\overline{L_{j}}$–derivatives and the constant in the $\mathcal{O}(\|u\|^{2})$ is independent of $\eta$. We rewrite (4.2) as
(4.3) $\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\partial\overline{\partial}(-h_{\eta})\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)\lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}\\ -2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\big((T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})u_{jK}\big)\overline{u_{kK}}+\mathcal{O}(\left\|u\right\|^{2})\;.$
@@
(4.4) $\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(\tilde{L}_{u}^{K})\right|^{2}\lesssim(1-\eta)\Big(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}\big((T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})u_{jK}\big)\overline{u_{kK}}+\mathcal{O}(\left\|u\right\|^{2})\Big)\\ +M_{\eta}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}u_{jK}\overline{u_{kK}}.$
@@
(4.5) $\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\left|\alpha_{\eta}(\tilde{L}_{u}^{K})\right|^{2}\lesssim(1-\eta)\bigg(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}\\ -2\text{Re}\big(\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{jk}\int_{b\Omega}c_{jk}\big((T-\tilde{M}_{\eta})u_{jK}\big)\overline{u_{kK}}\,\big)+\mathcal{O}(\left\|u\right\|^{2})\bigg)$
@@
It now suffices to prove that $\sum^{{}^{\prime}}_{|J|=q-1}\int_{b\Omega}|\alpha_{\eta}(\tilde{L}_{\mathcal{P}^{k,\eta}u}^{J})|^{2}$ is dominated by the right hand side of (4.1), for $k=+,0,$-.
@@
(4.21) $\sideset{}{{}^{\prime}}{\sum}_{J}|\alpha_{\eta}(\tilde{L}^{J}_{\mathcal{P}^{+,\eta}u})|^{2}\\ \lesssim(1-\eta)\left(\|\overline{\partial}_{b}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b,-h_{\eta}}^{*}(e^{-h_{\eta}/2}\mathcal{P}^{+,\eta}u)\|_{-h_{\eta}}^{2}+\|\overline{\partial}_{b}u\|^{2}+\|\overline{\partial}_{b}^{*}u\|^{2}\right)\\ +C_{\eta}\|u\|_{-1}^{2}\;.\;\;\;\;\;\;\;\;$
@@
(4.22) $2\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\partial\overline{\partial}(-h_{\eta})\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)\\ +2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}(T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})+3q\left\|\overline{L}u\right\|^{2}\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\mathcal{O}(\left\|u\right\|^{2})+3q\left\|\overline{L}u\right\|^{2}.$
@@
(4.24) $2\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\partial\overline{\partial}(-h_{\eta})\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)+2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}(T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})u_{jK}\overline{u_{kK}}\\ -3q\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}(c_{jj}Tu_{J})\overline{u_{J}}+3q\left\|Lu\right\|^{2}+\mathcal{O}(\left\|{L}u\right\|\left\|u\right\|)+\mathcal{O}(\left\|\overline{L}u\right\|\left\|u\right\|)\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\mathcal{O}(\left\|u\right\|^{2})+3\left\|\overline{L}u\right\|^{2}.$
Using $\|\overline{L}u\|^{2}\geq 0$, the Cauchy-Schwarz inequality, maximal estimates, and (2.2) gives
(4.25) $2\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\partial\overline{\partial}(-h_{\eta})\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)\\ +2\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}(T+\frac{1}{2}\frac{\partial h_{\eta}}{\partial\nu})u_{jK}\overline{u_{kK}}-3q\sum_{j}\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}(c_{jj}Tu_{J})\overline{u_{J}}\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}.$
@@
(4.26) $2\sideset{}{{}^{\prime}}{\sum}_{K}\int_{b\Omega}\partial\overline{\partial}(-h_{\eta})\big(\tilde{L}_{u}^{K}\wedge\overline{\tilde{L}_{u}^{K}}\big)\\ +\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\left(\left(2c_{jk}-3\delta_{jk}\text{Tr}(c)\right)T+\frac{\partial h_{\eta}}{\partial\nu}c_{jk}\right)u_{jK}\overline{u_{kK}}\\ \lesssim\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}\;,$
where $Tr(c)$ denotes the trace of the Levi form. Multiplying both sides of (4.26) by $(1-\eta)$, then adding $M_{\eta}\sum^{{}^{\prime}}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}u_{jK}\overline{u_{kK}}$, and invoking (3.8) and (2.2) leads to
(4.27) $\sideset{}{{}^{\prime}}{\sum}_{J}\int_{b\Omega}|\alpha_{\eta}(\tilde{L}_{u}^{J})|^{2}\\ \lesssim(1-\eta)\bigg(\|\bar{\partial}_{b}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\|\bar{\partial}_{b,-h_{\eta}}^{\star}(e^{-h_{\eta}/2}u)\|_{-h_{\eta}}^{2}+\left\|\bar{\partial}_{b}u\right\|^{2}+\left\|\bar{\partial}_{b}^{\star}u\right\|^{2}\\ -\text{Re}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}\left(2c_{jk}-3\delta_{jk}\text{Tr}(c)\right)Tu_{jK}\overline{u_{kK}}+\tilde{M}_{\eta}\sideset{}{{}^{\prime}}{\sum}_{K}\sum_{j,k}\int_{b\Omega}c_{jk}u_{jK}\overline{u_{kK}}\bigg),$
where $\tilde{M}_{\eta}$ is chosen big enough so that $\frac{M_{\eta}}{1-\eta}-\frac{\partial h_{\eta}}{\partial\nu}\leq\tilde{M}_{\eta}$. We now need an upper bound on the last two terms on the right hand side of (4.27), with $u$ replaced by $\mathcal{P}^{-,\eta}u$, equivalently, a lower bound on